# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ V-C2 Multi-Edge Collaboration
Multi-edge collaboration [240, 241, 242] provisions all data flow within the edge networks, where edge devices individually perform local computation on the local data they possess and collectively train or infer GI models by sharing information and merging the results. On the one hand, compared to the cloud-assisted solutions, multi-edge collaboration fully reserves data without leaking them to the remote datacenter, thus avoiding users’ privacy concerns from cloud providers [243, 112, 109]. On the other hand, collaborating multiple edge devices augments the available resources for targeted GI tasks beyond a single edge device, which therefore enables superior performance against on-device computation. Besides, privacy requirements are also guaranteed if the participating edge devices are mutually trusted, which can be relevant for many edge scenarios, e.g., in a smart home with edge devices owned by the same user and in a manufacturing pipeline with cameras and sensors possessed by the same company [244, 183, 245]. Under this paradigm, Fograph [246, 247] proposes a distributed GNN inference system that leverages multiple edge servers in proximity to IoT data sources for real-time GI model serving. To address resource heterogeneity and dynamics among edge servers, Fograph designs a load-aware inference execution planner as well as an agile workload scheduler for maximum parallelization. It also introduces a degree-aware compression mechanism to minimize data transmission overhead. Following this principle, GLAD [248] studies the cost optimization problem for distributed GNN serving in multi-tier edge networks. By formulating cost factors in the whole lifecycle of edge-enabled GI services, it separately considers static graphs and dynamics graphs and derives a holistic graph layout scheduling solution with theoretical performance guarantees.
VI Graph Intelligence for Edge Networks
@@
Besides directly applying GNN models to specific edge applications discussed in Sec. IV, GI techniques are also employed for optimizing edge networks. This can be relevant since many components in edge networks, e.g., network hierarchy as illustrated in Fig. 7(e), can be naturally recognized in graphs. Towards that, we discuss GI-based optimization on edge networks with respect to their resource management, graph-based decision-making, security issues, and GI-assisted FL in the following.
VI-A Network Resource Management
@@
Edge networks are booming with heterogeneous devices becoming part of them, and thus raise soaring demand for efficient network resource management. Given the graph structure of massively connected edge networks, GI models are deemed to be a useful tool to achieve that and are leveraged for optimizing network resource allocation and network function orchestration.
VI-A1 Resource Allocation
@@
Resource allocation is critical for ensuring networking efficiency and robustness, and has been extensively studied with traditional optimization means. Instead of applying complicated formulation and optimization, GI-based methods model the whole network as a graph and schedule an allocation strategy in a holistic manner [249].
@@
Power control in wireless networks seeks the optimal power allocation to transmitters to minimize power consumption while maintaining signal quality [250, 251]. Shen et al. [252] introduce an interference GCN to solve the power control problem in wireless networks, designed to model K-user interference channels as complete graphs and incorporating wireless channel data as graph features for scalability. Their approach is complemented by Eisen et al. [253, 254], who employed spectral-based GCNs for power allocation in Device-to-Device (D2D) wireless networks, highlighting the effectiveness of GCNs in managing interference. Building on these ideas, Nikoloska et al. [255] utilized REGNN to tackle power control challenges in decentralized wireless networks, leveraging the adaptability of GNNs to changing network topologies. Similarly, Yang et al. [256] propose UWGNN, which enhances GNNs by unrolling modules in the weighted minimum mean-square error algorithm into GI frameworks, demonstrating a method to integrate knowledge more effectively. Together, these studies underscore the versatility of GNNs in addressing power control problems across various wireless network settings, from centralized to decentralized architectures, and from static to dynamic topologies.
@@
Link scheduling is another key issue in communication networks as it drastically impacts the overall performance of the system (throughput, delay, etc.) [251]. To address the high computational complexity of traditional optimization methods, Lee et al. [257] proposed a graph embedding-based method for link scheduling in D2D networks. This approach effectively circumvents the need for accurate channel state information, making it more practical for real-world applications. In a similar vein, GCN-GAN [258] integrates the strengths of GCN, LSTM, and GAN to capture the dynamics, topology, and evolutionary patterns of weighted dynamic networks for temporal link prediction. These methods collectively leverage graph-based models to enhance link scheduling and prediction, demonstrating the versatility and effectiveness of graph neural networks in addressing various aspects of network performance optimization.
@@
VI-A2 Network Function Orchestration
@@
Virtual network function (VNF) virtualizes the functions that traditionally run on dedicated hardware devices such as routers and firewalls, and has become a key enabler in software-defined networks (SDNs) [262]. Researchers employ GI models on VNF for orchestrating various tasks including adaptive service function chain (SFC) and network slicing. SFC, crucial for identifying efficient paths in network servers for processing requested VNFs, faces challenges in maintaining high QoS in complex scenarios. Heo et al. [263] introduce a GNN-based architecture with an encoder-decoder model for dynamic network topologies, focusing on maintaining high QoS. The encoder represents the network topology, while the decoder evaluates nodes for processing VNFs, adapting to topology changes. Heo et al.’s subsequent work [264] overcomes the limitation of earlier approaches restricted to fixed topologies, by employing reinforcement learning for GNN-based model training across various unlabeled network topologies. Network slicing, as a promising paradigm in 5G networks, offers tailored resources for diverse services but faces challenges from virtualized infrastructure and various QoS requirements [265]. Pioneering studies [266, 267] employ digital twin (DT) technology to digitally replicate slicing-enabled networks, creating detailed virtual models that mirror the physical network’s dynamic characteristics. These models are then used to develop GNN algorithms that learn and optimize network behaviors directly from these non-Euclidean graph representations, thereby addressing challenges in resource allocation and QoS management in highly dynamic network environments.
VI-B Graph-Based Decision Making
@@
Given the powerful ability to abstract graphs, GI models can be integrated with DRL techniques to enable decision-making on graphs. Their confluence leads to Deep Graph Reinforcement Learning (DGRL), which is extensively utilized for making scheduling decisions on edge networks, e.g., offloading, routing, and wireless network control. In general, the DGRL framework can be illustrated in Fig. 12, where an agent observes the state, schedules configurations, and feeds the agent (neural network) back with a reward after each action. The GI model is incorporated as a graph information encoder to abstract relational semantics from the state.
VI-B1 Offloading and Routing
@@
Offloading and routing are classic techniques in edge networks that allow workload migration across networking entities and some researchers use graphs to abstract tasks to be scheduled [268, 269, 270] for offloading decisions. Specifically, ACED [268] introduces an actor-critic mechanism for DAGs-based computation offloading decisions in edge computing. By combining DRL with GCN, it optimizes offloading strategies, capturing application structures and task features, and considering task dependencies and wireless interference. This method significantly reduces the average energy-time cost for users by integrating topology and channel interference into decision-making, demonstrating the potential of GNNs in optimizing complex computation offloading scenarios. Swaminathan et al. [269] enhance a routing algorithm for SDN using a GNN functioning as a deep Q-Network. Trained to forecast efficient routing paths, this GNN minimizes packet delivery delays through simultaneous training and inference, enabling continuous learning and real-time adaptation. This approach illustrates the adaptability of GNNs in dynamic network environments, specifically focusing on optimizing routing decisions in SDN.
@@
VI-B2 Wireless-Enabled Control Systems
@@
DRL techniques have been frequently used in attaining efficient control of communication resources in wireless networks. However, trivially applying DRL does not scale well with the network size, and GI models are thus employed to overcome this challenge. Lima et al. [271] introduce a method that integrates DRL and GNNs for scalable, model-free resource allocation in large-scale wireless control systems. This method capitalizes on the ability of GNNs to apply spatial-temporal convolutions to graph-structured data, harnessing the natural graph form of wireless networks for efficient policy design. In their subsequent work [272], Lima et al. extend this approach by incorporating long-term constraints and demonstrate the permutation invariance and transferability of graph-based reinforcement learning policies in wireless control systems.
VI-C Security and Anomaly Detection
@@
Detecting anomalies in a graph is an important task in the analysis of complex systems, especially for ubiquitous edge networks. The anomalies, in the context of graphs, are patterns that fail to conform to normal patterns expected of the graph’s attributes or structures. In this respect, the community has developed GI models for anomaly detection tasks and applied to optimize edge networks.
VI-C1 Static Network
@@
For edge networks that are static within a period, e.g., a smart home network in a short slot, we call it a static network. Extensive research has been conducted focusing on the detection of anomalies within static graphs, encompassing various categories including vertex, link, and graph-level anomalies [273]. AnomMAN [274] is a GCN-based framework for anomalous node detection in multi-view attributed networks. It innovatively incorporates a graph auto-encoder module to leverage the low-pass filtering characteristic of graph convolution, typically a disadvantage for anomaly detection, transforming it into a beneficial feature by utilizing reconstruction errors to effectively identify anomalies. EFraudCom [275] introduces a fraud detection system employing competitive CGNN to identify anomalous edges in e-commerce platforms. Evaluations on real-world datasets demonstrate eFraudCom’s superior performance in detecting fraudulent activities. The system’s core innovation lies in modeling the distributions of normal and fraudulent behaviors separately, enhancing robustness against evolving fraud patterns. GLocalKD [276] and OCGTL [277] both tackle graph-level anomaly detection. The former introduces a method using random distillation of graph and node representations for deep anomaly detection, and the latter employs self-supervised and transformation learning to enhance existing deep one-class approaches [278].
VI-C2 Dynamic Network
@@
Dynamic edge networks take temporality into consideration where their topology and attributes may change over time [279]. Recent research in graph anomaly detection requires simultaneous modeling for time-series and graphs when both representations are present. Recent studies, such as those by Deng et al. [280] and Zhang et al. [281], have delved into the intricacies of multivariate time series anomaly detection within cyber-physical systems. Both works underscore the significance of understanding the complex interplay between sensors over time. Deng et al. [280] introduce an attention-based graph deviation network to learn the complex relationships between sensors in multivariate time series, not only accurately detecting anomalies but also providing an explainable model that identifies and clarifies deviations from normal sensor interactions. Similarly, Zhang et al. [281] introduce a method focusing on sensor dependence relationships in cyber-physical systems, utilizing a VAE for feature extraction and a GNN with stochastic learning to capture inter-sensor dependencies. Building upon these methodologies, FuSAGNet [282] combines Sparse Autoencoder with GNN for anomaly detection in high-dimensional time series, which optimizes both reconstruction and forecasting while explicitly modeling the relationships within multivariate time series.
VI-D GI-Assisted Federated Learning

### v2

@@ VI-B2 Multi-Edge Collaboration
Multi-edge collaboration [203, 204, 205] provisions all data flow within the edge networks, where edge devices individually perform local computation on the local data they possess and collectively train or infer GI models by sharing information and merging the results. On the one hand, compared to the cloud-assisted solutions, multi-edge collaboration fully reserves data without leaking them to the remote datacenter, thus avoiding users’ privacy concerns from cloud providers [206, 91, 207]. On the other hand, collaborating multiple edge devices augments the available resources for targeted GI tasks beyond a single edge device, which therefore enables superior performance against on-device computation [88, 208]. Besides, privacy requirements are also guaranteed if the participating edge devices are mutually trusted, which can be relevant for many edge scenarios, e.g., in a smart home with edge devices owned by the same user and in a manufacturing pipeline with cameras and sensors possessed by the same company [171, 209].
Under this paradigm, Fograph [210, 211] proposes a distributed GNN inference system that leverages multiple edge servers in proximity to IoT data sources for real-time GI model serving. To address resource heterogeneity and dynamics among edge servers, Fograph designs a load-aware inference execution planner as well as an agile workload scheduler for maximum parallelization. It also introduces a degree-aware compression mechanism to minimize data transmission overhead. Following this principle, GLAD [212] studies the cost optimization problem for distributed GNN serving in multi-tier edge networks. By formulating cost factors in the whole lifecycle of edge-enabled GI services, it separately considers static graphs and dynamics graphs and derives a holistic graph layout scheduling solution with theoretical performance guarantees. Both Fograph and GLAD are Level-4 systems for their functionality of GNN inference at the edge.
VI-C On-Device Computation
VI-C1 Resource-Efficient Procedure
Without modifying GI models, resource-efficient procedure design indicates the endeavors that develop GI processing flows in minimal resource usage, e.g., memory footprint. For instance, centering on the out-of-memory issues where GI models’ sizes are too large to fit into edge devices’ memory, Zhou et al. [213] (Level-3 rating) design a feature decomposition approach that decomposes the dimension of feature vectors and performs aggregation respectively, reducing up to 5$\times$ memory footprint against generic GI frameworks. Building on the theme of memory efficiency, SUGAR [103] (Level-4 rating) targets resource-efficient training on resource-constrained edge devices through a subgraph-level training scheme that first partitions the initial graph into a set of disjoint subgraphs and then performs local training for individual subgraphs. Complementing these approaches, RAIN [214] (Level-3 rating) explores the opportunity of conservatively processing similar inference batches based on local sensitive hash and reusing repeated vertices’ data among successive batches to reduce redundant data loading.
There are also some works that utilize channel pruning methods to reduce input feature dimensions for GI model inference acceleration. Zhou et al. [215] set the stage by employing a LASSO regression-based pruning technique, which identifies the most influential feature channels, thus ensuring that the pruning process is both effective and minimally disruptive to the model’s accuracy. Yik et al. [216] extend this line of inquiry by conducting a meticulous analysis of the trade-offs inherent in channel pruning. They develop algorithms that not only preserve the model’s predictive prowess but also enhance computational efficiency, aligning seamlessly with the work [215] by further honing in on the most valuable features.
VI-C2 Resource-Efficient Model
To automate the resource-efficient model design for edge platforms, some researchers [104, 220, 221] introduce NAS techniques to the GI domain and present hardware-aware NAS frameworks tailored to GNNs, aiming to strike a balance between model accuracy and efficiency corresponding to the characteristics of targeted devices. Zhou et al. [104] present HGNAS, a hardware-aware graph neural architecture search (GNAS) framework that optimizes GNN designs for edge devices by balancing accuracy and efficiency using a hardware performance predictor. Similarly, PaSca [220] constructs scalable GNNs through a novel scalable GNAS paradigm, enabling systematic exploration of a vast design space to optimize GNNs for both accuracy and efficiency via multi-objective optimization. Building upon these concepts, MaGNAS [221] is a mapping-aware GNAS framework that efficiently processes vision-based GNN workloads on heterogeneous multi-processor platforms by identifying optimal GNN architectures and mappings for maximized resource efficiency and performance trade-offs. Together, these works highlight the growing trend towards hardware-aware and scalable GNAS frameworks that address both the accuracy and efficiency requirements crucial for deploying GNNs on edge devices. This line of work represents Level-4 EGI systems since they enable both edge training and inference.
VI-C3 Model Compression
Compressing a complete model into a smaller one that appropriately fits the capability of the targeted edge device has been a popular deployment means for many edge intelligence services. The smaller model, which usually with fewer parameters, allows much less computing cost and thus significant acceleration for on-device computation. To obtain such a lightweight model for EGI, researchers typically employ three ways, i.e., model quantization, knowledge distillation, and neural architecture search.
Model quantization intends to quantize GI model parameters from an origin, larger bitwidth to a targeted, smaller bitwidth, and thus reduce the total size of the model. In this context, SGQuant [222] proposes a multi-granularity method that can quantize feature vectors at the component level, topology level, and layer level to meet diverse data precision demands. It also designs an automatic bitwidth-selecting algorithm to attain proper configuration among different quantization granularities. Building upon the quantization paradigm, Degree-Quant [223] pioneers a quantization-aware training scheme specifically crafted for GI models. This scheme is designed to facilitate low-precision inference on devices with constrained resources, without compromising the model’s predictive capabilities. VQ-GNN [224] takes a quantum leap in this domain by introducing a unified framework that applies vector quantization (VQ) for dimensionality reduction and learns a small number of quantized reference vectors of global node representations, which viably avoids the “neighbor explosion” problem and enables faster mini-batch training and inference of GI models. Following this quantization principle, some researchers go one step further with GI model binarization, i.e., binarizing parameters in GI models for aggressive execution speedup [225].
Knowledge Distillation (KD) aims at extracting knowledge from a complex teacher model and transferring it into a compact student model while maintaining comparable performance [226, 227], as depicted in Fig. 13 While KD is originally applied in the computer vision domain [228, 229, 230], Yang et al. [231] first introduce KD for GI models along with a local structure preserving method for teacher-student similarity measurement and experimentally show its effectiveness in GI-based 3D object detection tasks. Along with this line, Yang et al. [232] combine parameterized label propagation and feature-based MLP in the KD’s student model, allowing better utilization of the knowledge from both the teacher model and prior basement. TinyGNN [233] focuses the KD principle on accelerating model inference and designs a peer-aware module and a neighbor distillation strategy for preserving model accuracy. They all focus on the edge inference aspect and are Level-3 EGI systems in the rating taxonomy.
Figure 13: In general, KD for GNN follows a similar procedure of KD for traditional DL models, where we share the knowledge of the teacher model with the student model to improve its accuracy.
Lessons Learned (Sec. VI) Edge networks as a physical infrastructure can support GI in various architectures, including FGL with edge-cloud synergy, collaborative computing with distributed edge platforms, and in-situ computing on individual edge devices. The architecture used for EGI services depends on the demand of SLO (e.g., latency and privacy requirements) and the availability of the cloud. The lower reliance on the cloud the system has, the more stringent the constraint of computing resources, and thus the more attention is paid to resource-efficient optimizations.
VII Graph Intelligence for Edge Networks
@@
Besides directly using GI models to specific edge applications discussed in Sec. V, GI techniques are also employed for optimizing edge networks. This can be relevant since many components in edge networks, e.g., network hierarchy as illustrated in Fig. 7(e), can be naturally recognized in graphs. Towards that, we discuss GI-based optimization on edge networks with respect to their resource management, graph-based decision-making, security issues, and GI-assisted FL in the following. Using GI for optimizing edge not only involves GI model inference at the edge but also training tailored GI at targeted edge platforms, thus their EGI systems are at Level 4 in the rating taxonomy.
VII-A Network Resource Management
@@
Given the graph structure of massively connected edge networks, GI models are deemed to be a useful tool to achieve that and are leveraged for optimizing network resource allocation and network function orchestration.
VII-A1 Resource Allocation
@@
Resource allocation is critical for ensuring networking efficiency and robustness, and has been extensively studied using traditional optimization means. Instead of applying complicated formulation and optimization, GI-based methods model the whole network as a graph and schedule an allocation strategy in a holistic manner [234, 235, 236].
@@
Power control in wireless networks seeks the optimal power allocation to transmitters to minimize power consumption while maintaining signal quality [237, 238]. Shen et al. [239] introduce an interference GCN to solve the power control problem in wireless networks, designed to model K-user interference channels as complete graphs and incorporate wireless channel data as graph features for scalability. Their approach is complemented by Eisen et al. [240, 241], who employed spectral-based GCNs for power allocation in Device-to-Device (D2D) wireless networks, highlighting the effectiveness of GCNs in managing interference. Salaün [242] study cell-free MIMO networks and formulate a graph of the dominant dependence relationship between access points and user equipment, using GNN for scheduling the power control. Building on these ideas, Nikoloska et al. [243] utilized REGNN to tackle power control challenges in decentralized wireless networks, leveraging the adaptability of GNNs to changing network topologies. Similarly, Yang et al. [244] propose UWGNN, which enhances GNNs by unrolling modules in the weighted minimum mean-square error algorithm into GI frameworks, demonstrating a method to integrate knowledge more effectively. Together, these studies underscore the versatility of GNNs in addressing power control problems across various wireless network settings, from centralized to decentralized architectures, and from static to dynamic topologies.
@@
Link and channel scheduling is another key issue in communication networks as it drastically impacts the overall performance of the system (throughput, delay, etc.) [238]. Some researchers apply GNN or combine GNN with other models for channel tracking and throughput optimization in MIMO networks, where the channel states and links are formulated in graphs [245, 242, 246, 247]. Yang et al. [245] intend to represent the obtained channel data in graphs by associating channel correlations as links’ weights, based on which a GNN model is trained for channel tracking. D2D networks are also studied in a similar vein, e.g., GCN-GAN [248] integrates the strengths of GCN, LSTM, and GAN to capture the dynamics, topology, and evolutionary patterns of weighted dynamic networks for temporal link prediction. These methods collectively leverage graph-based models to enhance link scheduling and prediction, demonstrating the versatility and effectiveness of graph neural networks in addressing various aspects of network performance optimization.
@@
VII-A2 Network Function Orchestration
Virtual Network Function (VNF) virtualizes the functions that traditionally run on dedicated hardware devices such as routers and firewalls and has become a key enabler in software-defined networks (SDNs) [252]. Researchers employ GI models on VNF for orchestrating various tasks including adaptive Service Function Chain (SFC) and network slicing. SFC, crucial for identifying efficient paths in network servers for processing requested VNFs, faces challenges in maintaining high QoS in complex scenarios.
@@ VII-A2 Network Function Orchestration
Heo et al. [253] introduce a GAE for dynamic network topologies, where the encoder represents the network topology and the decoder evaluates nodes for processing VNFs, adapting to topology changes. Heo et al.’s subsequent work [254] overcomes the limitation of earlier approaches restricted to fixed topologies, by employing RL for GNN-based model training across various unlabeled network topologies. Pioneering studies on network slicing [255, 256] employ Digital Twin (DT) technology to digitally replicate slicing-enabled networks, creating detailed virtual models that mirror the physical network’s dynamic characteristics. These models are then used to develop GNNs that learn and optimize network behaviors directly from these non-Euclidean graph representations, thereby addressing challenges in resource allocation and QoS management in highly dynamic network environments. Besides the above aspects, we refer readers interested more about GNN for communications to related discussions [31, 257].
VII-B Graph-Based Decision Making
@@
Given the powerful ability to abstract graphs, GI models can be integrated with DRL techniques to enable decision-making on graphs. Their confluence leads to DGRL, which is extensively utilized for scheduling decisions on edge networks, e.g., offloading, routing, and wireless network control. In general, the DGRL framework can be illustrated in Fig. 14, where an agent observes the state, schedules configurations, and feeds the agent (neural network) back with a reward after each action. The GI model is incorporated as a graph information encoder to abstract relational semantics from the state.
VII-B1 Offloading and Routing
@@
Offloading and routing are classic techniques in edge networks that allow workload migration across networking entities and some researchers [101, 102, 258] use graphs to abstract tasks to be scheduled for offloading decisions. Specifically, ACED [101] introduces an actor-critic mechanism for DAGs-based computation offloading decisions considering task dependencies and wireless interference. Swaminathan et al. [102] enhance a routing algorithm for SDN using a GNN functioning with a deep Q-Network. Trained to forecast efficient routing paths, this GI model minimizes packet delivery delays through simultaneous training and inference, enabling continuous learning and real-time adaptation.
@@
VII-B2 Wireless-Enabled Control Systems
@@
DRL techniques have been frequently used in attaining efficient control of communication resources in wireless networks. However, trivially applying DRL does not scale well with the network size, and GI models are thus employed to overcome this challenge. For example, Nakashima et al. [259] design a DGRL method for scalable, model-free resource allocation in large-scale wireless control systems. It capitalizes on the ability of GNNs to apply spatial-temporal convolutions to graph-structured data, harnessing the natural graph form of wireless networks for efficient policy design. Some of its follow-up works [260, 261] extend this approach by incorporating long-term constraints and demonstrate the permutation invariance and transferability of graph-based reinforcement learning policies in wireless control systems.
VII-C Security and Anomaly Detection
@@
Detecting anomalies in a graph is an important task in the analysis of complex systems, especially for ubiquitous edge networks. The anomalies, in the context of graphs, are patterns that fail to conform to normal patterns expected of the graph’s attributes or structures. In this respect, the community has developed GI models for anomaly detection tasks and applied them to optimize edge networks. Fig. 15 illustrates its rationale, where vertices (with features) can be transformed into embeddings, which are used for anomaly identification in a latent space.
VII-C1 Static Network
@@
For edge networks that are static within a period, e.g., a smart home network in a short slot, we call it a static network. Extensive research has been conducted focusing on the detection of anomalies within static graphs, encompassing various categories including vertex, link, and graph-level anomalies [262]. AnomMAN [263] incorporates a graph auto-encoder module to leverage the low-pass filtering characteristic of graph convolution, typically a disadvantage for anomaly detection, transforming it into a beneficial feature by utilizing reconstruction errors to effectively identify anomalies. EFraudCom [264] proposes a fraud detection system employing GI to identify anomalous edges in e-commerce platforms. The system’s core innovation lies in modeling the distributions of normal and fraudulent behaviors separately, enhancing robustness against evolving fraud patterns. GLocalKD [265] and OCGTL [266] both tackle graph-level anomaly detection. The former introduces a method using random distillation of graph and node representations for deep anomaly detection, and the latter employs self-supervised and transformation learning to enhance existing deep one-class approaches [267].
Figure 15: GI models can abstract vertices’s features into embeddings, which can be used for detecting anomalies by classifying in a latent space.
VII-C2 Dynamic Network
@@
Dynamic edge networks take temporality into consideration where their topology and attributes may change over time [268]. Recent studies, such as those by Deng et al. [269] and Zhang et al. [270], have delved into the intricacies of multivariate time series anomaly detection within cyber-physical systems. Both of them underscore the significance of understanding the complex interplay between sensors over time. Deng et al. [269] propose an attention-based graph deviation network to learn the complex relationships between sensors in multivariate time series, not only accurately detecting anomalies but also providing an explainable model that identifies and clarifies deviations from normal sensor interactions. Similarly, Zhang et al. [270] focus on sensor dependence relationships in cyber-physical systems, utilizing a VAE for feature extraction and a GNN with stochastic learning to capture inter-sensor dependencies. Building upon these methodologies, FuSAGNet [271] combines Sparse Autoencoder with GNN for anomaly detection in high-dimensional time series, which optimizes both reconstruction and forecasting while explicitly modeling the relationships within multivariate time series.
VII-D GI-Assisted Federated Learning